# Methods to Estimate Cryptic Sequence Complexity

### v1

@@ Introduction
Understanding under what conditions and through what mechanisms complexity evolves is a significant open question in artificial life and evolutionary biology (taylor2016open; pigliucci2009extended). Indeed, numerous dimensions of complexity in biological systems have been considered, including those of developmental processes, phenotypic traits, and ecological interactions (szathmary2001can; mcshea2000functional). Among dimensions of biological complexity, the information content of genetic sequences is often useful due to its relative tractability, generalizability across systems, and foundational role to other aspects of biological complexity (adami2002complexity). Although information theoretic formulations have been established to describe biological sequence complexity (weiss2000information), counts of adaptive sites (i.e., genome sites that benefit fitness) are a convenient, commonly-used proxy measure for genome sequence complexity (dolson2019modes).
However, challenges can arise in identifying adaptive sites within a genome. Beyond very small genome sizes, complete identification of adaptive sites is hindered by combinatoric effects that make all-combinations analyses necessary to fully detangle epistatic effects effectively intractable (nitash2021information; adami2000evolution). This is particularly the case for systems with implicit fitness conditions with extensive biotic selection effects (moreno2022exploring; channon2000towards). In such circumstances, it can become necessary to use head-to-head competition trials between wildtype strains and knockout variants in situ to detect fitness effects (moreno2021case). Such competition-based fitness assays typically have sensitivity limitations, which limit detection of sites with small fitness effects. Here, we term genome sites with adaptive effects that are not directly detectable through single-site knockouts as “cryptic” sequence complexity.
To enable more complete sequence complexity analyses of digital organisms inclusive of “cryptic” adaptive sites, we propose three assays to conduct statical estimates of
@@
    Additive effect cryptic sites: sites with small contributions to fitness that, individually, fall below the threshold of detectability of fitness assays,
@@
    Epistatic effect cryptic sites: sites with fitness effects that are only observable in the context of other knockouts (i.e., redundancy),
@@
    Any effect cryptic sites: sites with any contribution to fitness, inclusive of the above.
The following sections describe proposed assays and report initial experiments with a simple model system designed to validate their estimations of cryptic sequence complexity. Software and data, including full, documented implementations of proposed underlying statistical estimators, are available via GitHub at https://github.com/mmore500/cryptic-sequence-concept.
Additive Effect Sites
@@
The assay proceeds by sampling sets of remaining sites are sampled and knocking them out together. In the presence of small-effect sites, a classic dose-response curve will occur as knockout set size is increased. That is, detectable fitness effects will be observed more frequently for larger knockout sets. The shape of this dose curve will depend on both the abundance of small-effect sites and their mean effect size. Fitting a negative binomial distribution to the curve allows these factors to be estimated. This distribution models, for coin flips with success probability $p$, the number of successive trials required to achieve $n$ successes. If we consider sampled sites being small-effect versus true-neutral as a coin flip event, then the proportion of small-effect sites in the genome will correspond to $p$ and the number of small-effect mutations required to reach detectability, $n$, will be inverse to the mean per-site effect size. Under this framing, the dose-response curve corresponds to the cumulative distribution function of an underlying negative binomial distribution.
@@
For an initial experiment exploring this approach, we generated a simple genome with 1,000 sites, 50 of which were designated to be small-effect sites. Effect sizes for these sites were uniformly between 0 and 0.2, relative to an effect detectability threshold of 1.0. To decide dose levels for the main assay, we first sampled 250 doses spread evenly from 1 to 1,000 sites and tested for detectable fitness effects from one sample at each dose level. We used the interval between the lowest dose with a detected fitness effect and the highest dose without a detected fitness effect as the dosing range for the main assay, choosing five dosing levels spaced evenly across this range. In this case, selected dose levels were 87, 136, 186, 236, and 286 sites. We then performed 1,000 knockouts of site sets sampled at each dosage then fit a negative binomial distribution to the observed outcome frequencies. Resulting estimates of the additive site count and mean effect size were both accurate, matching the true values of 50.0 and 0.1, respectively.
Epistatic Effect Sites
Figure 1: Use of knockout effect sizes and inclusion rates within “skeletonized” minimal viable genomes to distinguished small-effect and epistatic genome sites.
@@
This assay is designed to detect sites that only express detectable fitness knockout effects in the presence of specific other knockouts due to redundant masking. For this assay we generated minimal fitness-equivalent genome “skeletons” by knocking out sites until no more could be removed without detectable fitness loss. Because only one among a set of redundant sites will appear in any skeleton, the frequency with which redundant sites are excluded from skeletons should be high. However, some very small effect additive sites will also be able to be eliminated from skeletons. To disambiguate these scenarios, we perofrm an additional step: “jackknife” one-by-one knockouts of each site within sampled skeletons. Jackknifed sites with detectable, but very small magnitude fitness effects can then be identified as likely additive, rather than epistatic.
@@
For an initial experiment exploring this approach, we generated a sample genome with 4,000 sites, 200 of which were designated to be small-effect sites. Effect sizes for these sites were uniformly between 0 and 0.7, relative to the detectability threshold of 1.0. We additionally introduced 20 sets of 5 redundant sites. Fitness penalties uniformly distributed between 0.7 and 1.6 units were incurred when all sites within a set were knocked out.
To perform the assay, we first generated 20 sample skeletons. After performing Jackknife knockouts of each skeleton, we plotted skeleton exclusion frequency versus mean jackknife severity for each site that appeared in any skeleton. As shown in Figure 1, epistatic sites fall into the upper right quadrant. This procedure was able to identify 80 of the true 94 epistatic sites in the sample genome.
Any Effect Sites
@@
Any genome site with some fitness benefit should, in principle, potentially appear within a genome skeleton. Put another way, skeletons sample randomly (though not uniformly) from genome sites that provide some fitness benefit. The composition of genome skeletons can thus be analogized to the content of traps used by wildlife biologists to estimate the size of animal populations. Robust statistical procedures have been developed to account for biasing factors, including variation in capture probabilities among individuals (“trap shyness”) (amstrup2010handbook). In this work, we use the Burnham-Overton procedure to estimate the total number of any-effect sites from the distribution of capture counts among sites appearing within skeleton genomes (burnham1979robust).

### v2

@@ Introduction
Understanding under what conditions and through what mechanisms complexity evolves is a significant open question in artificial life and evolutionary biology (taylor2016open; pigliucci2009extended). Indeed, numerous dimensions of complexity in biological systems have been considered, including those of developmental processes, phenotypic traits, and ecological interactions (szathmary2001can; mcshea2000functional). Among dimensions of biological complexity, the information content of genetic sequences is often useful due to its relative tractability, generalizability across systems, and foundational role to other aspects of biological complexity (adami2002complexity). Although information theoretic formulations have been established to describe biological sequence complexity (weiss2000information), counts of adaptive sites (i.e., genome sites that benefit fitness) are a convenient, commonly-used proxy measure for genome sequence complexity (dolson2019modes; moreno2021case).
However, challenges can arise in identifying adaptive sites within a genome. Beyond very small genome sizes, complete identification of adaptive sites is hindered by combinatoric effects that make all-combinations analyses necessary to fully detangle epistatic effects effectively intractable (nitash2021information; adami2000evolution). This is particularly the case for systems with implicit fitness conditions with extensive biotic selection effects (moreno2019toward; channon2000towards). In such circumstances, it can become necessary to use head-to-head competition trials between wildtype strains and knockout variants in situ to detect fitness effects (moreno2022exploring). Such competition-based fitness assays typically have sensitivity limitations, which limit detection of sites with small fitness effects. Here, we term genome sites with adaptive effects that are not directly detectable through single-site knockouts as “cryptic” sequence complexity.
To enable more complete sequence complexity analyses of digital organisms inclusive of “cryptic” adaptive sites, we propose three assays to conduct statistical estimates of
@@
    Additive-effect cryptic sites: make small contributions to fitness that, individually, fall below the threshold of detectability of fitness assays;
@@
    Epistatic-effect cryptic sites: exhibit fitness effects only in the context of other knockouts (i.e., redundancy); and
@@
    Any-effect cryptic sites: make any contribution to fitness, inclusive of the above.
The following sections describe proposed assays and report initial experiments with a simple model system designed to validate their estimations of cryptic sequence complexity. Software, including full, documented implementations of underlying statistical estimators, is at https://github.com/mmore500/cryptic-sequence-concept (moreno2024cryptic).
Additive-effect Cryptic Sites
@@
The assay proceeds by sampling sets of remaining sites are sampled and knocking them out together. In the presence of small-effect sites, a classic dose-response curve will occur as knockout set size is increased. That is, detectable fitness effects will be observed more frequently for larger knockout sets. The shape of this dose-response curve will depend on both the abundance of small-effect sites and their mean effect size. Fitting a negative binomial distribution to the curve allows these factors to be estimated. This distribution models, for coin flips with success probability $p$, the number of successive trials required to achieve $n$ successes. If we consider sampled sites being small-effect versus true-neutral as a coin flip event, then the proportion of small-effect sites in the genome will correspond to $p$ and the number of small-effect mutations required to reach detectability, $n$, will be inverse to the mean per-site effect size. Under this framing, the dose-response curve corresponds to the cumulative distribution function of an underlying negative binomial distribution.
@@
For an initial experiment exploring this approach, we generated a simple genome with 1,000 sites. Fifty sites were designated to be small-effect sites, with effect sizes uniformly distributed between 0 and 0.7 (relative to a detectability threshold of 1.0). To decide dose levels for the main assay (i.e., numbers of sites in knockout samples), we first sampled 250 doses spread evenly from 1 to 1,000 sites and tested for detectable fitness effects from one sample at each dose level. We used the interval between the lowest dose with a detected fitness effect (25 sites) and the highest dose without a detected fitness effect (140 sites) as the dosing range for the main assay, choosing five dosing levels spaced evenly across this range. We performed 1,000 knockouts of site sets sampled at each dosage level and then fit a negative binomial distribution to the observed outcome frequencies. Resulting estimates of the additive site count and mean effect size were 30.0 and 0.5, approximating the true values of 50 sites and 0.35 effect size.
Epistatic-effect Cryptic Sites
Figure 1: Distinguishing between small-effect and epistatic genome sites. Epistatic sites exhibit both 1) severe fitness effects when knocked out individually from “skeletonized” minimal viable genomes (i.e., “jackknifed”) and 2) are often absent from sampled “skeleton” genomes (i.e., high exclusion rates). Minor jitter added to points for clarity.
@@
This assay is designed to detect sites that only express detectable fitness knockout effects in the presence of specific other knockouts due to redundant masking. For this assay we generated minimal fitness-equivalent genome “skeletons” by knocking out sites until no more could be removed without detectable fitness loss. Because only one among a set of redundant sites will appear in any skeleton, the frequency with which redundant sites are excluded from skeletons should be high. However, some very small effect additive sites will also be able to be eliminated from skeletons. To disambiguate these scenarios, we perform an additional step: “jackknife” one-by-one knockouts of each site within sampled skeletons. Jackknifed sites with detectable, but very small magnitude fitness effects can then be identified as likely additive, rather than epistatic.
@@
For an initial experiment exploring this approach, we generated a test genome with 4,000 sites, 200 of which were designated to be small-effect sites. Effect sizes for these sites were uniformly distributed between 0 and 0.7, relative to the detectability threshold of 1.0. We additionally introduced 20 sets of 5 redundant sites. Fitness penalties uniformly distributed between 0.7 and 1.6 units were incurred when all sites within a set were knocked out.
To perform the assay, we first sampled 20 genome skeletons. We then compared the frequency with which each site appeared in any skeleton to the severity of fitness effects when knocked out individually. As Figure 1 shows, epistatic sites are distinct in having both high skeleton exclusion rates and high jackknife knockout severity. Applied to the sample genome, this procedure successfully identified 80 of the true 94 epistatic sites.
Any-effect Cryptic Sites
@@
Any genome site with some fitness benefit should, in principle, potentially appear within a genome skeleton. Put another way, skeletons sample randomly (though not uniformly) from genome sites that provide some fitness benefit. The composition of genome skeletons can thus be analogized to the content of traps used by wildlife biologists to estimate the size of animal populations. Robust statistical procedures exist to account for biasing factors, including variation in capture probabilities among individuals (“trap shyness”) (amstrup2010handbook). We use the Burnham-Overton procedure to estimate the total number of any-effect sites from the distribution of capture counts among skeleton genome sites (burnham1979robust).